OpenStudy (anonymous):

1.solve for x (a)abslutevalue of(4x-1)+absol(x-2)=4 (b)Absol(11x-22)=-11

OpenStudy (hba):

|4x-1| + |x-2|=4

OpenStudy (openstudier):

\[\left| 4x-1 \right|+\left| x-2 \right|=4\]

OpenStudy (anonymous):

Yeah it is like this thanks guys!!!

OpenStudy (hba):

|x-2|=4-|4x-1|

OpenStudy (anonymous):

i tried but i nedd ur answer

OpenStudy (hba):

What answer did you get ?

OpenStudy (hba):

@neba.zi ?

OpenStudy (anonymous):

solution set are x=-1/5,7/5,1

OpenStudy (hba):

Yeah now check.

OpenStudy (anonymous):

i tired using sign chart first u find th zero of 4x-1& x-2 then arrange increasing order x<1/4 1/4 (1/4,2) 2 x>2 4x-1 --------------0++++++++++++++++ x-2 -------------------------------0++++++ let us take cases Case1If X<1/4 -(4x-1)-(x-2)=4 solve this u will get x=-1/5 Case2 If 1/4<x<2 4x-1-(x-2)=4 then solve this x=1 Case3 x>2 4x-1+x-2=4 again solve this x=7/5 Finally u got three solutions that satisfy the These are -1/5,1,7/5 I think it is clear OK try b Hurry up!!!!!

OpenStudy (hba):

You got these three answer but you have to check by putting x=-1/5,1,7/5 And you will see that 7/5 is not a solution :) so your final answer would be x=-1/5,1

OpenStudy (anonymous):

I think the 2nd question is easy Hurry up!!! Take a great care for this question!!!!